# Turbocharging constraints on dark matter substructure through a synthesis of strong lensing flux ratios and extended lensed arcs

### v2

@@ 4.3.2 Modeling of lens and source light profiles
We model the main deflector light as an elliptical Sérsic profile. We model the quasar emission region as a circular Gaussian, and sample its size uniformly $\mathcal{U}\left(1,10\right)\ \rm{pc}$ for each system. We model the quasar host galaxy with an ellpitical Sérsic profile with additional small-scale complexity implemented through shapelet basis functions (Birrer et al., 2015). An integer $n_{\rm{max}}$ determines the degree of complexity these functions can describe, with the complexity of the reconstructed image with increasing $n_{\rm{max}}$.
@@
where $k$ is the number of model parameters, $n$ represents the number of data points, and $\max\left[\mathcal{L}\right]$ represents the maximum likelihood of the data given the model. The BIC statistic compensates between obtaining a better fit to data and over-fitting a model. To calculate the BIC, we fit a smooth lens model (i.e. a lens model without substructure) to each mock. For each mock system, we found the BIC reaches a minimum for a source model consisting of an elliptical Sérsic profile plus shapelets at $n_{\rm{max}}=10$. With this source model we obtain a reduced $\chi^{2}$ per degree of freedom $\chi^{2}/\rm{DOF}\approx 1.2$ for the lens models we accept based on the imaging data likelihood. Appendix A provides additional discussion regarding the reconstruction of the source light in our simulations.
@@ 5 Results
This section presents the results of the inference of the lens model and substructure properties obtained from 25 simulated strong lens systems analyzed using the inference methodology discussed in Section 2, the lens modeling approach discussed in Section 3, and the simulation details presented in Section 4. We begin in Section 5.1 with case studies of four mock lens systems. We discuss in detail how the image positions, flux ratios, and imaging data work in tandem to simultaneously constrain the mass profile of the main deflector and the properties of substructure in each system. In Section 5.2, we present the joint constraints on the normalization of the subhalo mass function and the free-streaming length of dark matter obtained from the full sample, and quantify the degree to which including constraints from lensed arcs improves constraints on substructure properties relative to existing analysis methods that use only image positions and flux ratios.
@@ 5.1 Case studies
We begin by analyzing the results of applying our inference methodology to four mock lenses to gain physical insight into how the joint modeling of lensed arcs, image positions and flux ratios constrains the lens model and the properties of substructure. Figures 5, 6, 7, and 8 show the four lenses chosen for these case studies. We pick these four systems because they exhibit a variety of image configurations, lens and source redshifts, experience varying degrees of perturbation by halos, and have among the most informative likelihood functions. The left panels show the simulated lensed images, while the right-hand panels display the true convergence in dark matter substructure for each mock. As in previous work (Gilman et al., 2019), we define the convergence for a multi-plane lens system in terms of the divergence of the deflection field (see Equation 3)
@@ 5.1.1 The complementary information conveyed by imaging data and flux ratios
As the imaging data subtends angular scales comparable to the image separation, we expect the reconstruction of the lensed arcs will impose the strongest constraints on the large-scale mass profile of the main deflector. Figures 9, 10, 11, and 12 show the inference on a subset of the parameters that describe the main deflector mass profile. The parameters shown in each figure include the normalization of the main deflector mass profile $\theta_{\rm{E}}$, the axis ratio $q$, the ellipticity position angle $\phi_{\rm{q}}$, the external shear strength $\gamma_{\rm{ext}}$, the external shear position angle $\phi_{\rm{ext}}$, and the amplitude of the $m=3$ and $m=4$ multipole moments $a_{3}$ and $a_{4}$. Contours show the $68\%$ and $95\%$ confidence intervals for the parameters after marginalizing over the properties of substructure in each lens. The black contours show constraints obtained from the ray tracing methods presented by Gilman et al. (2019) that use only image positions and flux ratios to constrain the lens model. Blue contours show constraints obtained from applying methods discussed in Sections 2.2 and 3 to compute the likelihood and model the image positions and imaging data. The magenta contours show constraints obtained from combining the flux ratio likelihood with the image position and imaging data likelihoods.
@@
The flux ratios, however, convey vital information for constraining the properties of the deflection field on angular scales below those probed by imaging data [^8] We recall that we have constructed the likelihood function in such a way that the posterior distribution of substructure properties given the data can only differ from the prior if one incorporates constraints from the flux ratios (see Section 2.2 and Equation 6).. Figure 13 shows the joint likelihood function computed with the image positions, flux ratios, and imaging data for each of the four case study lenses shown in Figures 5-8. The parameters shown in the figures are $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$, the normalization of the subhalo mass function (Equation 11) and the cutoff scale of the halo mass function (Equations 10 and 13). The color scale in Figure 13 corresponds to the relative likelihood between positions in the parameter space. The top-left region of parameter space includes dark matter models with a significant suppression of the halo mass function and relatively few subhalos, while the bottom right corner of parameter space includes a plethora of subhalos and a CDM-like halo mass function.
@@ 5.1.2 Visualizing accepted and rejected realizations
Figures 14-17 also serve to aid in the interpretation of the likelihood functions shown in Figure 13, with Mock $6$ providing a particularly clear illustration of how perturbations by dark matter halos manifest in the likelihood. Closely examining the true multiplane convergence map for Mock $6$ in right panel of Figure 6, a collection of dark matter subhalos and line-of-sight halos appear in close proximity to image A (the bottom-right image). The collective impact of these structures imparts a $\sim 12\%$ perturbation to the magnification of this image. To match this feature in the data, substructure models that match the flux ratios in Mock 6 tend to have a single massive halo, or a collection of low-mass halos, perturbing image A. These halos appear prominently in the convergence maps shown in the central panels of Figure 15. We see a similar pattern in the case of Mock $23$, in which accepted lens models tend to have substructure perturbing the magnification of image B.
@@
Figure 18: The posterior distribution for $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ obtained from modeling only image positions and flux ratios (black), and image positions, flux ratios, and imaging data (blue). These posteriors are marginalized over nuisance parameters that include the mass profile of the main deflector, including multipole moments $a_{3}$ and $a_{4}$, the finite-size of the quasar emission region, and the surface brightness of the lensed quasar host galaxy. Black and blue vertical bars in the marginal likelihood for $m_{\rm{hm}}$ correspond to $95\%$ exclusion limits. The red crosshairs indicates the ground-truth used to create the simulated data, and the vertical lines in the marginal likelihood for $m_{\rm{hm}}$ correspond to $95\%$ exclusion limits. For visualization we have marked the CDM ground truth $m_{\rm{hm}}=0$ as $m_{\rm{hm}}=10^{4.2}M_{\odot}$.
@@ 5.2 Inference on substructure properties from the full sample
Figure 18 shows the joint inference on the normalization of the subhalo mass function, $\Sigma_{\rm{sub}}$, and the half-mode mass, $m_{\rm{hm}}$, obtained from 25 mock lenses. The black contours show the inference resulting from applying the ray tracing methods presented by Gilman et al. (2019), which use only the image positions and flux ratios to constrain the lens model and substructure properties. The blue contour in Figure 18 shows the constraints resulting from applying the methodology presented in Section 2, in which we use the lensed arcs, image positions, and flux ratios to constrain the lens model. To make a direct comparison between these two approaches and evaluate the relative improvement, we compute the likelihood obtained from only the image positions and flux ratios (black contours) using the same tolerance threshold $\epsilon$ to down-select on the flux ratio summary statistics when computing the blue likelihood, which uses all of the available data.
In terms of confidence intervals, incorporating imaging data improves the $95\%$ exclusion level by 0.5 dex; with only image positions and flux ratios, we find $\log_{10}m_{\rm{hm}}<10^{7.7}M_{\odot}$, and $\log_{10}m_{\rm{hm}}<10^{7.2}M_{\odot}$ when incorporating imaging data with a log-uniform prior on $m_{\rm{hm}}\in\mathcal{U}\left(4,10\right)$. We can also quantify the improvement gained by incorporating imaging data in terms of relative likelihoods, which do not depend on the prior. First, we define a region of parameter space associated with CDM as having $4.54.0<\log_{10}m_{\rm{hm}}/M_{\odot}<$ 4.5, and regions of parameter space associated with WDM as bins in $\log_{10}m_{\rm{hm}}$ with a width of 0.5 dex between $10^{7}M_{\odot}$ and $10^{9}{M_{\odot}}$. We then compute the relative likelihood between CDM and WDM as the volume of the posterior with $\log_{10}m_{\rm{hm}}<{4.5}$ to the volume of the posterior with $\log_{10}m_{\rm{hm}}$ in each bin.
@@ 6 Conclusions
The methodology we present can be applied to any dark matter model that predicts the form of the halo mass function and halo density profiles. For example, in self-interacting dark matter theories halos can undergo core collapse, a process that significantly increases their central density and therefore their lensing efficiency (Gilman et al., 2021; Minor et al., 2021; Yang & Yu, 2021; Gilman et al., 2023; Nadler et al., 2023). In fuzzy dark matter models, wave interference effects produce density fluctuations in the lens mass profile that impact both flux ratios and lensed arcs (Laroche et al., 2022; Powell et al., 2023). In terms of a direct test of a key prediction of cold dark matter, incorporating constraints from lensed arcs increases sensitivity to perturbation by low-mass substructure, as indicated by the stronger constraints on mass function turnovers on scales below $10^{7.5}M_{\odot}$. This increased sensitivity will aid in pushing constraints from strong lensing to scales below the threshold of galaxy formation, to scales $\sim 10^{7}M_{\odot}$ and below.
@@
The number of mock lenses we have analyzed in this work is partially motivated by the number of strong lens systems for which we currently have archival HST imaging data of lensed arcs and flux ratio measurements suitable for a milli-lensing analysis of dark substructure. Suitable flux ratios for milli-lensing must come from a region surrounding the background quasar spatially extended by $\gtrsim 1\rm{pc}$ such that it becomes immune to contamination from stellar micro-lensing. Such measurements can come from observations of nuclear narrow-line emission from Keck and the HST (Nierenberg et al., 2014; Nierenberg et al., 2017, 2020), radio measurements from VLBI (Koopmans et al., 2004; McKean et al., 2007; Hsueh et al., 2020), or emission from the warm dust region measured with JWST (Nierenberg et al., 2023). The synthesis of strong lensing flux ratios and extended lensed arcs, in combination with these various datasets from ground and space-based observatories, advances the observational frontier of cosmic probes of dark matter physics to uncharted territory.
@@ Acknowledgments
DG thanks Jo Bovy for his continuous support throughout the course of this project. We also thank Tommaso Treu for useful discussions.
@@ Appendix A The imaging data likelihood
Several features apparent in Figure 19 dissuade us from using the imaging data likelihood to directly constrain the properties of substructure. First, we see that randomly generated populations of halos, i.e. the “wrong” substructure models, sometimes result in better fits to the imaging data than the “correct” population of halos when we have imperfect knowledge of the source. Second, models with fewer halos (shown in red) systematically improve the imaging data likelihood relative to models with many halos (black). Both of these features could arise from degeneracies between small-scale structure in the lens mass distribution and small-scale features in the source light. These considerations motivate the inclusion of the importance sampling weights in Equation 6, which one can interpret as an adjustment of the prior volume associated with the source light model such that all dark matter models are equally likely when constrained by imaging data alone.

### v5

@@ 4.3.2 Modeling of lens and source light profiles
We model the main deflector light as an elliptical Sérsic profile. We model the quasar emission region as a circular Gaussian, and sample its size uniformly from a prior $\mathcal{U}\left(1,10\right)\ \rm{pc}$ for each system. We model the quasar host galaxy with an ellpitical Sérsic profile with additional small-scale complexity implemented through shapelet basis functions (Birrer et al., 2015). An integer $n_{\rm{max}}$ determines the degree of complexity these functions can describe, with the complexity of the reconstructed image with increasing $n_{\rm{max}}$.
@@
where $k$ is the number of model parameters, $n$ represents the number of data points, and $\max\left[\mathcal{L}\right]$ represents the maximum likelihood of the data given the model. The BIC statistic compensates between obtaining a better fit to data and over-fitting a model. To calculate the BIC, we fit a smooth lens model (i.e. a lens model without substructure) to each mock. For each mock system, we found the BIC reaches a minimum for a source model consisting of an elliptical Sérsic profile plus shapelets at $n_{\rm{max}}=10$. With this source model we obtain a reduced $\chi^{2}$ per degree of freedom $\chi^{2}/\rm{DOF}\approx 1.2$, with DOF=3694, for the lens models we accept based on the imaging data likelihood. Appendix A provides additional discussion regarding the reconstruction of the source light in our simulations.
@@ 5 Results
This section presents the results of the inference of the lens model and substructure properties obtained from 25 simulated strong lens systems assuming a CDM ground truth, and 25 lens systems created assuming a WDM ground truth. Both sets of lens systems are analyzed using the inference methodology discussed in Section 2, the lens modeling approach discussed in Section 3, and the simulation details presented in Section 4. We begin in Section 5.1 with case studies of four mock lens systems created with a CDM ground truth. We discuss in detail how the image positions, flux ratios, and imaging data work in tandem to simultaneously constrain the mass profile of the main deflector and the properties of substructure in each system. In Section 5.2, we present the joint constraints on the normalization of the subhalo mass function and the free-streaming length of dark matter obtained from the full sample, and quantify the degree to which including constraints from lensed arcs improves constraints on substructure properties relative to existing analysis methods that use only image positions and flux ratios.
@@ 5.1 Case studies
We begin by analyzing the results of applying our inference methodology to four mock lenses created with a CDM ground truth to gain physical insight into how the joint modeling of lensed arcs, image positions and flux ratios constrains the lens model and the properties of substructure. Figures 5, 6, 7, and 8 show the four lenses chosen for these case studies. We pick these four systems because they exhibit a variety of image configurations, lens and source redshifts, experience varying degrees of perturbation by halos, and have among the most informative likelihood functions. The left panels show the simulated lensed images, while the right-hand panels display the true convergence in dark matter substructure for each mock. As in previous work (Gilman et al., 2019), we define the convergence for a multi-plane lens system in terms of the divergence of the deflection field (see Equation 3)
@@ 5.1.1 The complementary information conveyed by imaging data and flux ratios
As the imaging data subtends angular scales comparable to the image separation, we expect the reconstruction of the lensed arcs will impose the strongest constraints on the large-scale mass profile of the main deflector. Figures 9, 10, 11, and 12 show the inference on a subset of the parameters that describe the main deflector mass profile. The parameters shown in each figure include the normalization of the main deflector mass profile $\theta_{\rm{E}}$, the axis ratio $q$, the ellipticity position angle $\phi_{\rm{q}}$, the external shear strength $\gamma_{\rm{ext}}$, the external shear position angle $\phi_{\rm{ext}}$, and the amplitude of the $m=3$ and $m=4$ multipole moments $a_{3}$ and $a_{4}$. Contours show the $68\%$ and $95\%$ credible intervals for the parameters after marginalizing over the properties of substructure in each lens. The black contours show constraints obtained from the ray tracing methods presented by Gilman et al. (2019) that use only image positions and flux ratios to constrain the lens model. Blue contours show constraints obtained from applying methods discussed in Sections 2.2 and 3 to compute the likelihood and model the image positions and imaging data. The magenta contours show constraints obtained from combining the flux ratio likelihood with the image position and imaging data likelihoods.
@@
The flux ratios, however, convey vital information for constraining the properties of the deflection field on angular scales below those probed by imaging data [^8] We recall that we have constructed the likelihood function in such a way that the posterior distribution of substructure properties given the data can only differ from the prior if one incorporates constraints from the flux ratios (see Section 2.2 and Equation 2.2.3).. Figure 13 shows the joint likelihood function computed with the image positions, flux ratios, and imaging data for each of the four case study lenses shown in Figures 5-8. The parameters shown in the figures are $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$, the normalization of the subhalo mass function (Equation 11) and the cutoff scale of the halo mass function (Equations 10 and 13). The color scale in Figure 13 corresponds to the relative likelihood between positions in the parameter space. The top-left region of parameter space includes dark matter models with a significant suppression of the halo mass function and relatively few subhalos, while the bottom right corner of parameter space includes a plethora of subhalos and a CDM-like halo mass function.
@@ 5.1.2 Visualizing accepted and rejected realizations
Figures 14-17 also serve to aid in the interpretation of the likelihood functions shown in Figure 13, with Mock $6$ providing a particularly clear illustration of how perturbations by dark matter halos manifest in the likelihood. Closely examining the true multiplane convergence map for Mock $6$ in right panel of Figure 6, a collection of dark matter subhalos and line-of-sight halos appear in close proximity to image A (the bottom-right image). The collective impact of these structures imparts a $\sim 12\%$ perturbation to the magnification of this image. To match this feature in the data, substructure models that match the flux ratios in Mock 6 tend to have a single massive halo, or a collection of low-mass halos, perturbing image A. These halos appear prominently in the convergence maps shown in the central panels of Figure 15.
@@
Figure 18: The posterior distribution for $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ obtained from modeling only image positions and flux ratios (black), and image positions, flux ratios, and imaging data (magenta). These posteriors are marginalized over nuisance parameters that include the mass profile of the main deflector, including multipole moments $a_{3}$ and $a_{4}$, the finite-size of the quasar emission region, and the surface brightness of the lensed quasar host galaxy. Black and magenta vertical bars in the marginal likelihood for $m_{\rm{hm}}$ correspond to $95\%$ exclusion limits. The red crosshairs indicates the ground-truth used to create the simulated data, and the vertical lines in the marginal likelihood for $m_{\rm{hm}}$ correspond to $95\%$ exclusion limits. For visualization we have marked the CDM ground truth $m_{\rm{hm}}=0$ as $m_{\rm{hm}}=10^{4.2}M_{\odot}$. The inference assumes flux ratio measurement precision of 3 percent and astrometric precision of 0.005 milli-arcseconds.
@@
Figure 19: The joint posterior on the normalization of the subhalo mass function $\Sigma_{\rm{sub}}$ and the half-mode mass $m_{\rm{hm}}$ for a simulated dataset with a WDM ground truth $m_{\rm{hm}}=10^{7.5}M_{\odot}$. The posterior results from flux ratio measurement uncertainties of 1 percent. Black contours show the posterior with a log-uniform prior on $\Sigma_{\rm{sub}}$, and magenta contours show the result of incorporating a Gaussian prior on $\Sigma_{\rm{sub}}$ centered on the ground truth value of $\log_{10}\Sigma_{\rm{sub}}=-1.4$ with a width 0.2 dex. As in Figure 18, contours correspond to $68\%$ and $95\%$ credible intervals, the red crosshairs mark the input ground truth, and the vertical bars in the $m_{\rm{hm}}$ marginal likelihood represent $95\%$ exclusion limits.
Figure 20: The same as Figure 19, but assuming flux ratio measurement uncertainties of 3 percent.
@@ 5.2 Inference on substructure properties from the full sample
[…]
In terms of credible intervals, incorporating imaging data improves the $95\%$ exclusion level by 0.5 dex; with only image positions and flux ratios, we find $\log_{10}m_{\rm{hm}}<10^{7.7}M_{\odot}$, and $\log_{10}m_{\rm{hm}}<10^{7.2}M_{\odot}$ when incorporating imaging data with a log-uniform prior on $m_{\rm{hm}}\in\mathcal{U}\left(4,10\right)$. We can also quantify the improvement gained by incorporating imaging data in terms of relative likelihoods, which do not depend on the prior. First, we define a region of parameter space associated with CDM as having $4.54.0<\log_{10}m_{\rm{hm}}/M_{\odot}<$ 4.5, and regions of parameter space associated with WDM as bins in $\log_{10}m_{\rm{hm}}$ with a width of 0.5 dex between $10^{7}M_{\odot}$ and $10^{9}{M_{\odot}}$. We then compute the relative likelihood between CDM and WDM as the volume of the posterior with $\log_{10}m_{\rm{hm}}<{4.5}$ to the volume of the posterior with $\log_{10}m_{\rm{hm}}$ in each bin.
@@
[…]
Our inference method recovers the input values for these parameters, even after marginalizing over the main deflector mass profile including third and fourth order multipole perturbations, the size of the warm dust region surrounding the lensed quasar, and the lensed quasar host galaxy light. The covariance between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ manifests more prominently for these inferences than for the CDM ground truth (Figure 18), but incorporating an informative prior for the amplitude of the subhalo mass function can aid in breaking this covariance. In practice, such a prior could come from N-body simulations or semi-analytic models (e.g. Benson, 2012; Fiacconi et al., 2016; Jiang et al., 2021; Nadler et al., 2023a; Mansfield et al., 2024; Du et al., 2024), which make increasingly robust predictions for the number of main deflector subhalos that appear in projection near the Einstein radius in typical host halos $\sim 10^{13}M_{\odot}$. Alternatively, a prior on the amplitude of the subhalo mass function could come from measurements of the halo mass function in lenses from gravitational imaging (e.g. Vegetti et al., 2014; Hezaveh et al., 2016; He et al., 2022; Wagner-Carena et al., 2023). The marginal likelihood of $m_{\rm{hm}}$ excludes CDM for the posteriors that assume a prior on $\Sigma_{\rm{sub}}$, with $\log_{10}m_{\rm{hm}}/M_{\odot}>6.7$ and $\log_{10}m_{\rm{hm}}/M_{\odot}>5.2$ for flux ratio measurement uncertainties of $1\%$ and $3\%$, respectively.
Comparing Figures 19 and 20, the inference on WDM parameters is particularly sensitive to the measurement uncertainty in the flux ratios. The systems observed by JWST (Nierenberg et al., 2024) have typical uncertainties of $3\%$, but we have included the inference assuming $1\%$ uncertainties to demonstrate that our inference methodology can recover input ground truth with sufficient measurement precision. The constraints obtained assuming $1\%$ precision suggest that a larger sample of lenses with $3\%$ uncertainties could yield a statistically significant constrain on WDM models with a turnover $m_{\rm{hm}}\sim 3\times 10^{7}M_{\odot}$.
@@ 6 Conclusions
  • Our method can recover the free-streaming cutoff in a WDM mass function, although the ability to detect a WDM cutoff near $10^{7}M_{\odot}$ requires measurement precision of $1\%$, or a larger sample of lenses with both flux ratio measurements and lensed arcs than the 25 considered here. Theoretically-motivated priors for the amplitude of the subhalo mass function can aid in breaking covariance between the number of subhalos and the free-streaming cutoff.
@@
The methodology we present can be applied to any dark matter model that predicts the form of the halo mass function and halo density profiles. For example, in self-interacting dark matter theories halos can undergo core collapse, a process that significantly increases their central density and therefore their lensing efficiency (Gilman et al., 2021; Minor et al., 2021; Yang & Yu, 2021; Gilman et al., 2023; Nadler et al., 2023b). In fuzzy dark matter models, wave interference effects produce density fluctuations in the lens mass profile that impact both flux ratios and lensed arcs (Laroche et al., 2022; Powell et al., 2023). In terms of a direct test of a key prediction of cold dark matter, incorporating constraints from lensed arcs increases sensitivity to perturbation by low-mass substructure, as indicated by the stronger constraints on mass function turnovers on scales below $10^{7.5}M_{\odot}$. This increased sensitivity will aid in pushing constraints from strong lensing to scales below the threshold of galaxy formation, to scales $\sim 10^{7}M_{\odot}$ and below.
@@
The number of mock lenses we have analyzed in this work is partially motivated by the number of strong lens systems for which we currently have archival HST imaging data of lensed arcs and flux ratio measurements suitable for a milli-lensing analysis of dark substructure. Suitable flux ratios for milli-lensing must come from a region surrounding the background quasar spatially extended by $\gtrsim 1\rm{pc}$ such that it becomes immune to contamination from stellar micro-lensing. Such measurements can come from observations of nuclear narrow-line emission from Keck and the HST (Nierenberg et al., 2014; Nierenberg et al., 2017, 2020), radio measurements from VLBI (Koopmans et al., 2004; McKean et al., 2007; Hsueh et al., 2020), or emission from the warm dust region measured with JWST (Nierenberg et al., 2024). The synthesis of strong lensing flux ratios and extended lensed arcs, in combination with these various datasets from ground and space-based observatories, advances the observational frontier of cosmic probes of dark matter physics to uncharted territory.
@@ Acknowledgments
DG thanks Jo Bovy for his continuous support throughout the course of this project. We also thank Tommaso Treu for useful discussions. We also thank the anonymous referee for a very constructive report that improved the presentation of the methodology.
@@ Appendix A The imaging data likelihood
Several features apparent in Figure 21 dissuade us from using the imaging data likelihood to directly constrain the properties of substructure. First, we see that randomly generated populations of halos, i.e. the “wrong” substructure models, sometimes result in better fits to the imaging data than the “correct” population of halos when we have imperfect knowledge of the source. Second, models with fewer halos (shown in red) systematically improve the imaging data likelihood relative to models with many halos (black). Both of these features could arise from degeneracies between small-scale structure in the lens mass distribution and small-scale features in the source light. These considerations motivate the inclusion of the importance sampling weights in Equation 2.2.3, which one can interpret as an adjustment of the prior volume associated with the source light model such that all dark matter models are equally likely when constrained by imaging data alone.
@@
Appendix B The effect of multipoles on substructure constraints
Shortly after this work appeared, Cohen et al. (2024) (hereafter C24) claimed that the presence of multipole perturbations to galaxy density profiles – in particular, the $m=3$ and $m=4$ moments that we considered in this work – preclude inferences of substructure properties from quadruply-imaged quasars. C24 draw this conclusion from a series of lens modeling experiements in which they fit a model that includes only $m=3$ and $m=4$ multipole perturbations to the image positions and flux ratios of mock lenses perturbed by substructure. C24 speculate that a minor difference in the modeling of the position angle of the $m=4$ term, $\phi_{4}$, explains the discrepancy between their findings and the constraints on substructure properties we obtain from analyzing only the image positions and flux ratios of 25 mock lenses (shown by the black posterior in Figure 18); we fix $\phi_{4}$ to the position angle of the underlying elliptical power law profile, while C24 allow it to vary freely.
Using the methods discussed in this paper, we can investigate the claims by C24 regarding the effect of multipole perturbations in substructure inferences. To begin, we repeat the inference presented in Section 5 assigning the same degree of model flexibility to the multipole terms as advocated by C24. Specifically, we sample $a_{3}$, $a_{4}$, and $\phi_{3}$ from the priors described in Section 4.2, but now allow $\phi_{4}$ to vary freely between $-\pi/8$ and $\pi/8$. Figure 22 shows the result of analyzing the 25 mock lenses with a CDM ground truth using the more flexible prior on $\phi_{4}$. We emphasize that the mocks in our sample have non-zero amplitudes of $a_{3}$ and $a_{4}$, and thus our simulations do not have priors centered on the ground truth.
After marginalizing over $a_{3}$, $a_{4}$, $\phi_{3}$, and $\phi_{4}$, we obtain the black posterior distribution shown in Figure 22. The blue posterior distribution in Figure 22 includes importance sampling weights that enforce alignment between the $m=4$ term and the underlying elliptical power law profile to isolate the effect of a freely-varying $\phi_{4}$ angle on the constraints. Assigning the same degree of flexibility to the lens model as advocated by C24, we see no evidence for a systematic bias in the inferred model parameters, or a degeneracy between multipoles and halos that would preclude the constraints shown in Figure 22.
To understand these results in the context of the claims by C24, we note that for most problems there exists some other model besides the one under consideration that can fit the dataset. C24 consider a model in which only multipoles can resolve flux ratio anomalies, and compute the required properties of the multipole terms in this scenario. However, to conclude that models with substructure are indistinguishable from lens models that include only multipoles, one must actually calculate the likelihood function and demonstrate that the data cannot distinguish the models statistically when both dark matter halos and multipole terms are present in the lens model. C24 fundamentally cannot make statements regarding relative likelihoods, and therefore the constraining power of the data, because they do not include substructure in the model they use to analyze mock data.
Models including substructure are strongly preferred from the perspective of Bayesian model selection, both in our sample of mock lenses and the four cases considered by C24. For the 25 mocks in this paper and the 4 generated by C24 we generate 600,000 possible lens models that include both dark matter substructure and multipole perturbations, and 600,000 lens models that include only multipole perturbations. We draw samples of $a_{3}$, $a_{4}$, $\phi_{3}$, and $\phi_{4}$ from the same prior distributions for each calculation, with the amplitudes and orientations of these terms allowed to vary freely, and we assign the same priors for $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ as discussed in Section 4.1. For the 29 systems in consideration, we perform a Bayesian model comparison by evaluating the posterior odds given the image positions and flux ratios
$\frac{p\left(M_{1}|\boldsymbol{O_{\rm{img}}},\boldsymbol{O_{\rm{f}}}\right)}{p\left(M_{2}|\boldsymbol{O_{\rm{img}}},\boldsymbol{O_{\rm{f}}}\right)}=\frac{p\left(M_{1}\right)}{p\left(M_{2}\right)}\frac{\mathcal{L}\left(\boldsymbol{O_{\rm{img}}},\boldsymbol{O_{\rm{f}}}|M_{1}\right)}{\mathcal{L}\left(\boldsymbol{O_{\rm{img}}},\boldsymbol{O_{\rm{f}}}|M_{2}\right)}.$ (17)
Here, $M_{1}$, represents the model that includes both substructure and multipoles, and $M_{2}$ represents the model that only includes multipoles. The first term on the right represents the ratio of our prior beliefs regarding the probability of $M_{1}$ and $M_{2}$, and the second term is the Bayes factor. For this test, we will ignore the multiple lines of evidence pointing towards the existence of dark matter substructure and assume that both models are equally likely. We can approximate the Bayes factor from the number of accepted samples generated under the Approximate Bayesian Computing approach outlined in Section 2. We use a stringent acceptance threshold $\epsilon<0.01$ for these calculations to match the flux ratios precisely.
Figure 23 shows the distribution of the posterior odds for each mock lens, with numbers greater than one indicating that a model with substructure and multipoles is preferred relative to a model that only includes multipoles. The 25 mocks we generate with a CDM ground truth are represented by the black histogram, and the odds for the four cases considered C24 are marked with vertical bars. One in four of the mocks considered by C24 exhibits a notable Bayesian preference (odds > 2) for substructure in the lens model, compared with one in five of the mocks we generate. The four mocks considered by C24 have a joint Bayes factor (the product of the individual Bayes factors) of 26, and for the 25 mocks we consider with $a_{3}$, $a_{4}$, $\phi_{3}$, and $\phi_{4}$ allowed to vary freely, we obtain a joint Bayes factor of 1140. From Figure 23, we conclude that the mock lenses considered by C24, which they use to support the claim that quadruply-imaged quasars cannot constrain substructure properties, actually exhibit a Bayesian preference for substructure in the lens model consistent with what one expects in CDM.
The reason we can constrain substructure properties, even with multipoles included in the lens model with the same degree of flexibility as advocated by C24, is that models including substructure reproduce the observed data more frequently than lens models that include only multipoles. As a result, the “multipoles-only” explanation for the data is heavily disfavored from a Bayesian standpoint. Accounting for lens models that include both substructure and multipole terms requires the careful evaluation of the integral in Equation 2, as outlined in this paper and in previous work (Gilman et al., 2019, 2020). These tests emphasize the importance of a rigorous statistical treatment of the problem in order to validate statements regarding the constraining power of certain datasets over different models, and the importance of validating modeling assumptions through tests on realistic datasets.
Figure 22: The posterior distribution for $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ obtained from modeling only image positions and flux ratios for the 25 lenses created with a CDM ground truth. The black posterior shows the result of analyzing the mock data while allowing the orientation of the $m=4$ multipole perturbation to vary freely, and the blue posterior shows the effect of adding importance weights that enforce alignment with the position angle of the underlying elliptical power law profile.
Figure 23: The posterior odds ratios, or Bayes factors (Equation B1), for the twenty-five mock lenses in our sample (black histogram), and the four mock lenses considered by C24 (colored vertical bars). A number greater than one indicates a Bayesian preference for a model that includes dark matter substructure and multipole perturbations, while a number less than one indicates a model with only multipoles is preferred. The dashed grey vertical bar marks a Bayes factor of one.